# Step-DeepResearch Technical Report

### v1

@@ Abstract
As Large Language Models (LLMs) shift toward autonomous agents, Deep Research has emerged as a pivotal metric for assessing the core competitiveness of agents. However, existing works primarily focus on academic multi-hop search tasks with ground truth, such as BrowseComp, which often struggle to satisfy user demands for open-ended research tasks in real-world scenarios. Open-ended research not only demands robust retrieval capabilities but also challenges the agent’s comprehensive skills in latent intent recognition, long-horizon decision-making, multi-turn tool use, logical structuring, and cross-source verification. To address this, we introduce Step-DeepResearch, a cost-effective, end-to-end deep research agent model. We propose a novel Data Synthesis Strategy Based on Atomic Capabilities, aimed at reinforcing underlying capabilities in planning, information gathering, reflection, and report writing. In terms of the training paradigm, we construct a progressive path from agentic mid-training to supervised fine-tuning and reinforcement learning. Combined with a Checklist-style Judger reward design, this approach significantly improves robustness across diverse scenarios. Furthermore, to address the lack of evaluations reflecting real-world demands in the Chinese domain, we establish ADR-Bench, a Chinese benchmark for realistic deep research scenarios. Experimental results demonstrate that Step-DeepResearch, with only 32B parameters, achieves a high score of 61.4% on the Scale AI Research Rubrics. In expert human evaluations on ADR-Bench, its Elo score significantly outperforms comparable models and rivals state-of-the-art closed-source models such as OpenAI DeepResearch and Gemini DeepResearch. These findings prove that through a refined training scheme, medium-sized models can achieve expert-level deep research capabilities. With extremely low deployment and inference costs, Step-DeepResearch stands as the most cost-effective deep research agent model currently available in the industry.
@@ 1 Introduction
Guided by this principle, we introduce Step-DeepResearch, an end-to-end framework that aims to enable robust and practical autonomous research primarily through native model capabilities. Our contributions are:
@@
With 32B parameters, Step-DeepResearch shows strong performance relative to its scale. On ADR-Bench, it achieves high usability in end-to-end research. Unlike many deep research systems that rely on complex multi-agent coordination or heavyweight workflows, Step-DeepResearch benefits from internalized atomic capabilities, using only a streamlined ReAct-style single-agent design. In expert-based Elo ratings, it not only outperforms larger models like MiniMax-M2 [6], GLM-4.6 [7], and DeepSeek-V3.2 [8], but also surpasses cutting-edge deep research systems such as Kimi-Researcher [3] and MiniMax Agent [6]. On the Research Rubrics Benchmark [9], Step-DeepResearch achieved 61.4% rubric compliance under ternary grading, performing at a level comparable to OpenAI Deep Research [1] and Gemini Deep Research [2], and significantly outperforming a range of open-source and proprietary models. This work demonstrates that models with medium-scale parameters can also achieve expert-level deep research capabilities. As shown in Figure 1, with its low deployment and inference costs, Step-DeepResearch has become the most cost-effective deep research system.
@@ 2.3 Deep Research Related Evaluation Benchmarks
Another category of benchmarks focuses more on retrieval and knowledge testing capabilities. BrowseComp [4] contains 1,266 fact-finding questions requiring multi-hop reasoning, directly targeting the evaluation of models’ ability to retrieve deeply hidden information. HLE (Humanity’s Last Exam) [21] includes 2,500 expert-level questions across multiple disciplines, yet still carries a strong “closed-book exam” flavor, largely focusing on multi-hop search with definitive answers or frontier knowledge testing, failing to fully reach the openness and multidimensionality of real industrial-grade research scenarios. To this end, we construct ADR-Bench, aiming to fill the evaluation gap for Chinese deep research scenarios driven by real user demands.Claude is AI and can make mistakes. Please double-check responses.
@@ 3 Data Strategy: Constructing Atomic Capabilities
Direct exploration within the raw token-level space $\mathcal{A}_{\text{token}}$ is not only computationally expensive but also prone to trapping the model in local optima due to the exponential growth of the branching factor with sequence length. Therefore, we proposes a unified data construction perspective: reshaping the training objective from “predicting the next token” to “deciding the next Atomic Action.”
@@ 3.1 Capability I: Planning & Task Decomposition
Planning capability requires the model to effectively decompose ambiguous or broad user requests into executable sub-tasks and to dynamically adjust its route based on environmental feedback. To achieve this, we adopt a “Reverse Engineering” strategy, utilizing existing “perfect planning results” from the real world to synthesize high-complexity planning data. Furthermore, to prevent tasks from being too trivial, we implemented a rigorous screening process for task queries.
@@ 3.2.3 Difficulty Filtering
To ensure the non-triviality of planning tasks, we use QwQ-32b [24] as a difficulty filtering model. Any Query that the QwQ-32b model can solve under the default ReAct framework is considered a “simple problem” and is excluded from the training set. Since QwQ-32b shares the same base model and pre-training knowledge as the model used in this work, but lacks extensive agent-specific training, tasks solvable by QwQ-32b can be regarded as simple tasks that do not require specialized training.
@@ 3.3.1 The Error-Reflection Loop
The specific synthesis pipeline is as follows: First, an expert-level model generates a preliminary search and reasoning trajectory. If the final output matches the answer, the trajectory is included directly in the training set as a positive sample. If the result does not match, we construct a Prompt asking the model to perform a reflection action based on the incorrect result. Based on the reflection conclusion, the model retains historical memory and retries the trajectory generation; this process iterates up to 3 times.
@@ 3.3.2 Deep Verification Workflow
Generated verification trajectories undergo strict posterior filtering. We input the [Verification Point, Conclusion, Evidence] triplet into a Judger model to verify if the report conclusion is completely logically self-consistent with the evidence, ensuring the model learns a verification paradigm that is both factually accurate and logically rigorous.
@@ 3.4 Capability IV: Report Generation
Report writing is not merely text generation; it is a process of structured reorganization of collected fragmented information. We divide the training of this capability into a Mid-training phase, focusing on domain style and content depth, and an SFT phase, focusing on instruction following and formatting specifications.
@@ 4.1 Stage 1: Agentic Mid-Training
Figure 2: Performance trends during mid-training (32K context) across increasing training token scales. Checkpoints are saved approximately every 5 billion tokens. The figure reports the average performance on SimpleQA, TriviaQA, and FRAMES benchmarks. As the training progresses, consistent performance improvements are observed across all benchmarks, with particularly significant gains on FRAMES, indicating enhanced agent-related and structured reasoning capabilities. Notably, the curves have not yet fully converged at 150 billion tokens, suggesting further potential for improvement.
As shown in Fig. 2, during training, checkpoints are saved approximately every 5 billion tokens, and performance gains are analyzed over the full 150-billion-token mid-training process in Stage I. The results indicate that as the training data scale increases, the model exhibits consistent performance improvements on the SimpleQA, TriviaQA, and FRAMES benchmarks. The maximum observed gains reach approximately +1.26% on SimpleQA, +2.30% on TriviaQA, and +10.88% on FRAMES. Overall, the 32K mid-training stage has not yet fully converged at the 150-billion-token scale, suggesting substantial room for further improvement, particularly in agent-related and structured reasoning capabilities.
@@
Building upon the stable convergence of capabilities achieved in Stage I, the second stage of mid-training further extends the maximum context length to 128K, with a primary focus on real-world complex task scenarios. This stage strengthens the model’s abilities in retrieval, planning, and tool-augmented reasoning under ultra-long contexts. The training data in this stage is more closely aligned with practical applications and covers complex structures such as web interaction, search processes, and multi-tool collaboration. The detailed data composition is summarized in Table 2.
@@
4.2 Stage 2: Post-training Supervised Fine-tuning (SFT)
@@
  • Deep Research data. This category targets comprehensive research tasks involving open-ended questions. The data covers the full pipeline of intent understanding, planning, information cross-verification, and report generation under strict formatting requirements (e.g., citation and attribution). Through this design, the entire end-to-end logic of intent analysis–planning–execution–reflection–writing is reinforced.
@@
4.3 Stage 3: Reinforcement Learning (RL)
In the first two stages, the training process primarily relied on large-scale synthetic and distilled data for mid-training and supervised fine-tuning (SFT). Although this approach is effective for instruction-following tasks and basic tool usage, it mainly depends on “teacher” trajectories and cannot be improved through trial-and-error learning with real-world interactions. Therefore, we introduces reinforcement learning (RL), connecting the model directly to the real tool usage environment and optimizing it through interactions with the environment. Unlike short-term searches focused on entity matching, the quality of deep research reports depends on multiple dimensions, including task decomposition, planning, tool invocation strategy, evidence selection and validation, and the final report generation. A recent approach is to transform report quality evaluation into a set of rubric-based scores, which are then used as optimization signals [27]. This rubric-based reinforcement learning enhances the model’s abilities in active planning, reflection, and cross-source validation, overcoming the limitations of offline imitation and significantly improving both performance and user experience.
@@ 4.3.1 RL Data Synthesis
  • In the first step, guided by a small number of high-quality examples and templates, we use a powerful LLM to generate an initial task description (referred to as the hidden task summary) and concurrently generate a set of fine-grained rubrics. Each rubric is defined as an atomic standard containing evaluation dimensions and importance weights, ensuring that the description is specific, verifiable, and practically applicable in task assessment. Furthermore, each rubric is assigned a role indicating whether it is an explicit, implicit, or negative requirement, corresponding to whether it needs to be explicitly stated in the task or serves as a bonus or penalty item.
  • In the second step, we generate the target task (i.e., the actual user task query) based on the generated rubrics. During this process, we simultaneously re-assess the role of each rubric for the synthesized task and provide brief reasoning and attribution. If any discrepancies are found between the initial and final role assignment, the entire synthesis sample is discarded. This ensures that the synthesized task queries align with the rubric’s requirements, semantically support the key dimensions in the rubric, and cover both explicit and implicit task requirements.
@@
To ensure that the objectives and requirements of the task remain aligned during the reverse synthesis process, we implement an objective consistency verification step to check the consistency between the hidden task summary, rubrics, and the synthesized task. This process is performed by an independent judge model, which evaluates the match between the hidden task summary and the synthesized task, as well as the alignment of the rubrics with the task’s requirements. The judge model assesses the relevance of each rubric and ensures no contradictions with the task objectives. Finally, it provides an overall consistency score and decides whether the sample should be retained. This step ensures high-quality task samples and reliable supervision signals for subsequent reinforcement learning training.
@@ 4.3.2 Reward Design
In the deep research report scenario, relying directly on expensive large LLMs to assign fine-grained scores based on the full rubrics for each report would require dozens of inferences per sample to cover all evaluation dimensions. This approach incurs prohibitively high computational costs and latency during large-scale agentic RL training. Based on our observation, the quality of the rubrics is often more critical in ensuring that the LLM Judge’s annotations align with human experts, rather than the absolute performance of the judge model itself. Thus, we first use a strong model to generate scores and explanations on the constructed (Query, Rubrics, Report) triples, which then serve as supervisory signals for training the Rubrics Judge model. This model undergoes two training phase: supervised fine-tuning and reinforcement learning with verifiable rewards. The former aims to help the model learn the scoring logic and explanation style of the strong model, establishing foundational discriminative ability, while the latter further strengthens the consistency of the model output format and the robustness of the scoring logic. Ultimately, this Rubrics Judge serves as the primary reward provider in this stage, allowing us to support large-scale agentic RL training within an acceptable budget without compromising signal quality.
@@
In our initial attempt to evaluate reports based on rubrics, we categorized each rubric’s judgment into three classes: “fully satisfied”, “partially satisfied” and “not satisfied”, mapping them directly to ternary judgment $\{1,0.5,0\}$. However, we found that the agreement between different strong models (acting as LLM judges) and between LLM judges and human experts was lower in the “partially satisfied” category. This misalignment in the intermediate category may led to unstable reward signals, making it difficult to provide consistent optimization directions and weakening the model’s ability to adhere to constraints. To ensure more distinguishable reward signals, we convert the ternary judgments into binary signals based on the nature of the rubrics:
@@
  • For negative rubrics: We map “not satisfied” to 0, while “partially satisfied” and “fully satisfied” are unified as $1$. This ensures that any deviations from the desired outcome are penalized, helping the model to avoid undesirable behaviors.
@@ 4.3.3 Agentic Reinforcement Learning
To enable the agent to genuinely learn how to complete deep-research tasks under real-world constraints, we place it in a multi-tool, multimodal interactive environment during the reinforcement learning stage. Within this environment, the agent can freely alternate between natural-language generation and tool invocation. For cost control, we introduce a caching mechanism for high-expense web retrieval, reusing results for identical queries, and impose explicit budgets on the number of tool calls and total token consumption to ensure that training operates within a predictable resource envelope. Within this environment, each deep-research task is treated as a reinforcement learning episode. The agent first performs intent analysis and restates the user query, then produces a step-by-step plan, organizing tool invocations and information gathering around distinct subproblems. It subsequently filters, contrasts, and cross-validates evidence from multiple sources, and ultimately drafts a complete report. After the episode terminates, the Rubrics Judge evaluates the report against the predefined rubrics, producing multi-dimensional scores that are converted—via the strict reward mapping into reward signals used for policy updates. By iterating this process, the model continuously learns through trial-and-error in a real environment, progressively acquiring policy decisions that more closely match human expert preferences with respect to tool selection, invocation timing, and call ordering.

### v2

@@ Abstract
As Large Language Models (LLMs) shift toward autonomous agents, Deep Research has emerged as a pivotal metric for assessing the core competitiveness of agents. However, existing works primarily focus on academic multi-hop search tasks with ground truth, such as BrowseComp, which often struggle to satisfy user demands for open-ended research tasks in real-world scenarios. Open-ended research not only demands robust retrieval capabilities but also challenges the agent’s comprehensive skills in latent intent recognition, long-horizon decision-making, multi-turn tool use, logical structuring, and cross-source verification. To address this, we introduce Step-DeepResearch, a cost-effective, end-to-end Deep Research agent model. We propose a novel Data Synthesis Strategy Based on Atomic Capabilities, aimed at reinforcing underlying capabilities in planning, information seeking, reflection, and report writing. In terms of the training paradigm, we construct a progressive path from agentic mid-training to supervised fine-tuning and reinforcement learning. Combined with a Checklist-style Judger reward design, this approach significantly improves robustness across diverse scenarios. Furthermore, to address the lack of evaluations reflecting real-world demands in the Chinese domain, we establish ADR-Bench, a Chinese benchmark for realistic Deep Research scenarios. Experimental results demonstrate that Step-DeepResearch, with only 32B parameters, achieves a high score of 61.42 on the Scale AI ResearchRubrics. In expert human evaluations on ADR-Bench, its Elo score significantly outperforms comparable models and rivals state-of-the-art proprietary services such as OpenAI DeepResearch and Gemini DeepResearch. These findings prove that through a refined training scheme, medium-sized models can achieve expert-level Deep Research capabilities. With extremely low deployment and inference costs, Step-DeepResearch stands as the most cost-effective Deep Research agent model currently available in the industry.
@@ 1 Introduction
Guided by these principles, we introduce Step-DeepResearch, an end-to-end framework that aims to enable robust and practical autonomous research primarily through native model capabilities. Our contributions are:
@@
With 32B parameters, Step-DeepResearch shows strong performance relative to its scale. On ADR-Bench, it achieves high usability in end-to-end research. Unlike many Deep Research systems that rely on complex multi-agent coordination or heavyweight workflows, Step-DeepResearch benefits from internalized atomic capabilities, using only a streamlined ReAct-style single-agent design. In expert-based Elo ratings, it not only outperforms larger models like MiniMax-M2 [6], GLM-4.6 [7], and DeepSeek-V3.2 [8], but also surpasses cutting-edge Deep Research systems such as Kimi-Researcher [3] and MiniMax Agent Pro [6]. On the ResearchRubrics Benchmark [9], Step-DeepResearch achieves 61.42 rubric compliance under ternary grading, performing at a level comparable to OpenAI DeepResearch [1] and Gemini DeepResearch [2], and significantly outperforming a range of open-source and proprietary models. This work demonstrates that models with medium-scale parameters can also achieve expert-level Deep Research capabilities. As shown in Figure 1, with its low deployment and inference costs, Step-DeepResearch has become the most cost-effective Deep Research system.
@@ 2.3 Deep Research Related Evaluation Benchmarks
Another category of benchmarks focuses more on retrieval and knowledge testing capabilities. BrowseComp [4] contains 1,266 fact-finding questions requiring multi-hop reasoning, directly targeting the evaluation of models’ ability to retrieve deeply hidden information. HLE (Humanity’s Last Exam) [21] includes 2,500 expert-level questions across multiple disciplines, yet still carries a strong “closed-book exam” flavor, largely focusing on multi-hop search with definitive answers or frontier knowledge testing, failing to fully reach the openness and multidimensionality of real industrial-grade research scenarios. To this end, we construct ADR-Bench, aiming to fill the evaluation gap for Chinese Deep Research scenarios driven by real user demands.
@@ 3 Data Strategy: Constructing Atomic Capabilities
Direct exploration within the raw token-level space $\mathcal{A}_{\text{token}}$ is not only computationally expensive but also prone to trapping the model in local optima due to the exponential growth of the branching factor with sequence length. Therefore, we propose a unified data construction perspective: reshaping the training objective from “predicting the next token” to “deciding the next Atomic Action.”
@@ 3.1 Capability I: Planning & Task Decomposition
Planning capability requires the model to effectively decompose ambiguous or broad user requests into executable sub-tasks and to dynamically adjust its route based on environmental feedback. To achieve this, we adopt a “Reverse Engineering” strategy, utilizing existing “perfect planning results” from the real world to synthesize high-complexity planning data. Furthermore, to prevent tasks from being too trivial, we implement a rigorous screening process for task queries.
@@ 3.2.3 Difficulty Filtering
To ensure the non-triviality of planning tasks, we use QwQ-32b [24] as a difficulty filtering model. Any query that the QwQ-32b model can solve under the default ReAct framework is considered a “simple problem” and is excluded from the training set. Since QwQ-32b shares the same base model and pre-training knowledge as the model used in this work, while lacks extensive agent-specific training, tasks solvable by QwQ-32b can be regarded as simple tasks that do not require specialized training.
@@ 3.3.1 The Error-Reflection Loop
The specific synthesis pipeline is as follows. An expert-level model generates a preliminary search and reasoning trajectory. If the final output matches the answer, the trajectory is included directly in the training set as a positive sample. If the result does not match, we construct a prompt asking the model to perform a reflection action based on the incorrect result. Based on the reflection conclusion, the model retains historical memory and retries the trajectory generation; this process iterates up to 3 times.
@@ 3.3.2 Deep Verification Workflow
The generated verification trajectories undergo strict posterior filtering. The verification point, conclusion, and evidence are treated as a triplet and input into a judge model to verify whether the report conclusion is logically self-consistent with the evidence, ensuring that the model learns a verification paradigm that is both factually accurate and logically rigorous.
@@ 3.4 Capability IV: Report Generation
Report writing is not merely text generation, it is a process of structured reorganization of collected fragmented information. We divide the training of this capability into a mid-training phase, focusing on domain style and content depth, and an SFT phase, focusing on instruction following and formatting specifications.
@@ 4.1 Stage 1: Agentic Mid-training
Figure 2: Performance trends during mid-training with 32K context. Average accuracy on SimpleQA, TriviaQA, and FRAMES is reported at checkpoints saved every  5B tokens. Performance improves steadily with training token scale, with particularly large gains on FRAMES, indicating strengthened agentic and structured reasoning. Curves remain unconverged at 150B tokens, suggesting further headroom.
As shown in Figure 2, during training, checkpoints are saved approximately every 5 billion tokens, and performance gains are analyzed over the full 150-billion-token mid-training process in Stage 1. The results indicate that as the training data scale increases, the model exhibits consistent performance improvements on the SimpleQA, TriviaQA, and FRAMES benchmarks. The maximum observed gains reach approximately +1.26% on SimpleQA, +2.30% on TriviaQA, and +10.88% on FRAMES. Overall, the 32K mid-training stage has not yet fully converged at the 150-billion-token scale, suggesting substantial room for further improvement, particularly in agent-related and structured reasoning capabilities.
@@
Building upon the stable convergence of capabilities achieved in Stage 1, the second stage of mid-training further extends the maximum context length to 128K, with a primary focus on real-world complex task scenarios. This stage strengthens the model’s abilities in retrieval, planning, and tool-augmented reasoning under ultra-long contexts. The training data in this stage is more closely aligned with practical applications and covers complex structures such as web interaction, search processes, and multi-tool collaboration. The detailed data composition is summarized in Table 2.
@@
4.2 Stage 2: Post-training Supervised Fine-tuning
@@
  • Deep Research data. This category targets comprehensive research tasks involving open-ended questions. The data covers the full pipeline of intent understanding, planning, information cross-verification, and report generation under strict formatting requirements (e.g., citation and attribution). Through this design, the entire end-to-end logic of intent-analysis–planning–execution–reflection–writing is reinforced.
@@
4.3 Stage 3: Reinforcement Learning
In the first two stages, the training process primarily relies on large-scale synthetic and distilled data for Mid-training and Supervised Fine-tuning (SFT). Although this approach is effective for instruction-following tasks and basic tool usage, it mainly depends on “teacher” trajectories and cannot be improved through trial-and-error learning with real-world interactions. Therefore, we introduce reinforcement learning (RL), connecting the model directly to the real tool usage environment and optimizing it through interactions with the environment. Unlike short-term searches focused on entity matching, the quality of Deep Research reports depends on multiple dimensions, including task decomposition, planning, tool invocation strategy, evidence selection and validation, and the final report generation. A recent approach is to transform report quality evaluation into a set of rubric-based scores, which are then used as optimization signals [27]. This rubric-based reinforcement learning enhances the model’s abilities in active planning, reflection, and cross-source validation, overcoming the limitations of offline imitation and significantly improving both performance and user experience.
@@ 4.3.1 RL Data Synthesis
  • In the first step, guided by a small number of high-quality examples and templates, we use a powerful LLM to generate an initial task description (referred to as the hidden task summary) and concurrently generate a set of fine-grained rubrics. Each rubric is defined as an atomic standard containing evaluation dimensions and importance weights, ensuring that the standard is specific, verifiable, and practically applicable in task assessment. Furthermore, each rubric is assigned a role indicating whether it is an explicit, implicit, or negative requirement, corresponding to whether it needs to be explicitly stated in the task or serves as a bonus or penalty item.
  • In the second step, we synthesize the target task (i.e., the actual user task query) based on the synthesized rubrics. During this process, we simultaneously re-assess the role of each rubric for the synthesized task and provide brief reasoning and attribution. If any discrepancies are found between the initial and final role assignment, the entire synthesis sample is discarded. This ensures that the synthesized task queries align with the rubric’s requirements, semantically support the key dimensions in the rubric, and cover both explicit and implicit task requirements.
@@
To ensure that the objectives and requirements of the task remain aligned during the reverse synthesis process, we implement an objective consistency verification step to check the consistency between the hidden task summary, rubrics, and the synthesized task. This process is performed by an independent judge model, which evaluates the consistency between the hidden task summary and the synthesized task, and the alignment of the rubrics with the task requirements. The judge model assesses the relevance of each rubric and ensures no contradictions with the task objectives. Finally, it provides an overall consistency score and decides whether the sample should be retained. This step ensures high-quality task samples and reliable supervision signals for subsequent reinforcement learning training.
@@ 4.3.2 Reward Design
In the Deep Research report scenario, relying directly on large LLMs to assign fine-grained scores based on the full rubrics for each report would require dozens of inferences per sample to cover all evaluation dimensions. This approach incurs prohibitively high computational costs and latency during large-scale agentic RL training. Based on our observation, the quality of the rubrics is often more critical in ensuring that the LLM Judge’s annotations align with human experts, rather than the absolute performance of the judge model itself.
Thus, we first use a strong model to generate scores and explanations on the constructed $\langle\text{Query, Rubrics, Report}\rangle$ triples, which then serve as supervisory signals for training the rubrics judge model. This model undergoes two training phases: supervised fine-tuning and reinforcement learning with verifiable rewards. The former aims to help the model learn the scoring logic and explanation style of the strong model, establishing foundational discriminative ability, while the latter further strengthens the consistency of the model output format and the robustness of the scoring logic. Ultimately, this Rubrics Judge serves as the primary reward provider in this stage, allowing us to support large-scale agentic RL training within an acceptable budget without compromising signal quality.
@@
In our initial attempt to evaluate reports based on rubrics, we categorized each rubric’s judgment into three classes: “fully satisfied”, “partially satisfied” and “not satisfied”, mapping them directly to ternary judgments $\{1,0.5,0\}$. However, we found that the agreement between different strong models (acting as LLM judges) and between LLM judges and human experts was lower in the “partially satisfied” category. This misalignment in the intermediate category may lead to unstable reward signals, making it difficult to provide consistent optimization directions and weakening the model’s ability to adhere to constraints. To ensure more distinguishable reward signals, we convert the ternary judgments into binary signals based on the nature of the rubrics:
@@
  • For negative rubrics: We map “not satisfied” to 0, while “partially satisfied” and “fully satisfied” are unified as $1$. This ensures that any deviations from the desired outcome are penalized, helping the model avoid undesirable behaviors.
@@ 4.3.3 Agentic Reinforcement Learning
To enable the agent to genuinely learn how to complete deep-research tasks under real-world constraints, we place it in a multi-tool interactive environment during the reinforcement learning stage. Within this environment, the agent can freely alternate between natural-language generation and tool invocation. For cost control, we introduce a caching mechanism for high-expense web retrieval, reusing results for identical queries, and impose explicit budgets on the number of tool calls and total token consumption to ensure that training operates within a predictable resource envelope. Within this environment, each deep-research task is treated as a reinforcement learning episode. The agent first performs intent analysis and restates the user query, then produces a step-by-step plan, organizing tool invocations and information gathering around distinct subproblems. It subsequently filters, contrasts, and cross-validates evidence from multiple sources, and ultimately drafts a complete report. After the episode terminates, the Rubrics Judge evaluates the report against the predefined rubrics, producing multi-dimensional scores that are converted—via the strict reward mapping into reward signals used for policy updates. By iterating this process, the model continuously learns through trial-and-error in a real environment, progressively acquiring policy decisions that more closely match human expert preferences with respect to tool selection, invocation timing, and call ordering.